#1 when table sugar (sucrose), C12H22O11(s) is digested, it is broken down by reacting with water to produce two simple sugars, glucose and fructose, each with the formula C6H12O6(s).

C12H22O11(s) + H2O(l) -----> C6H12O6(s) + C6H12O6(s)

An experiment was performed, and the following data were obtained ;( 4marks)

Run initial (sucrose) (mol/L) Rate (mol/(Ls))
1 0.050 3.09x10^-5
2 0.100 6.17x10^-5
3 0.250 1.50x10^-4

#2 For the reaction A + B +2C --> ABC2, the following data were obtained:

run (a) (mol/L) (B) (Mol/L) (C) (mol/L)
Rate (mol/Ls))
1 0.120 o.220 1.25 2.14x10^-3
2 0.120 0.220 2.50 2.14x10^-3
3 0.120 0.440 1.25 4.28x10^-3
4 0.240 0.220 1.25 8.56x10^-3

A) Determine the value of the rate law ( 4 marks )

B) Determine the rate law for this reaction ( 4 marks )
Answer #1

1) C12H22O11(s) + H2O(l) -----> C6H12O6(s) + C6H12O6(s)

Let the rate law is given by, Rate = k [C12H22O11]m k =rate constant

Given that, Rate =3.09x10-5mol/ls , [C12H22O11]= 0.050 mol/L -------------- (1)

Rate =6.17 x 10-5mol/ls , [C12H22O11]= 0.100 mol/L -------------- (2)

equation (2) / equation (1) ,(applying rate law)

(6.17 x 10-5) / (3.09x10-5) = (0.100 / 0.050)m

or, 2 = 2m

or, m = 1 i.e.- order w.r.t. C12H22O11 is 1 i.e.- 1st order

Thus the rate law is, Rate =k[C12H22O11]

let, Rate =3.09x10-5mol/ls , [C12H22O11]= 0.050 mol/L

k = Rate / [C12H22O11] = 3.09x10-5 mol/ls / 0.050 mol/L

or, k = 6.18 x 10-4 s-1